Kelly brought 8 apples, 3 pears, and 13 oranges and brought them home in a grocery bag. If she picks one from the bag without looking what is the probability of choosing an apple? a pear? an orange?
The chances of choosing an apple are________?
The chances of choosing an pear are________?
The chances oof choosing an orangeare________?

Answer by CubeyThePenguin(3113)

total amount of fruits = 8 + 3 + 13 = 24

apple: 8/24 = 1/3
pear: 3/24 = 1/8
orange: 13/24